OpenStudy (anonymous):

I really need help with this quastion a jet leaves an airport traveling at a steady rate of 600km another jet leaves the same airport 3/4 later traveling at 800km in the same direction. how long will it take the second jet to overtake the first

OpenStudy (icecube):

3/4? i m assuming it is 3/4hrs let v is the velocity first jet = 600km/hr and u is the velocity of second jet =800km/hr when second jet leave the airport the first is already covered a distance AB =600*3/4 AB= 450 relative velocity of second jet with respect first is V = v-u=200km/hr the time taken by second to overtake first is =AB/V T=450/200 =9/4hrs

OpenStudy (aroub):

d=r*t d=800t => first equation d-450=600t=> second equation 800t-450=600t 200t=450 t=9/4 d=800=9/4 d=1800 km So, it will take 2 hours and 15 min to take over the first

OpenStudy (radar):

Good job aroub, with clear steps.

OpenStudy (aroub):

Oh wait this d=800=9/4 is suppose to be d=800(9/4) Hehe, thank you radar!! =DD

OpenStudy (radar):

Yes, but the key is both planes will have traveled the same distance. The problem was only asking how long would the faster plane caught up with slower one and 9/4 hr (assuming the speeds were per hour)

OpenStudy (radar):

As I said "good answer"

OpenStudy (aroub):

extra information =)

OpenStudy (aroub):

No wait

OpenStudy (aroub):

so the distance is wrong?

OpenStudy (radar):

The distance would of been the fast plane speed times 9/4 (wouldn't it)

OpenStudy (icecube):

aroud urs distance and time both are correct

OpenStudy (icecube):

*aroub